OpenStudy (anonymous):

To buy both a new car and a new house, Tina sought two loans totaling $318,338. The simple interest rate on the first loan was 2.4%, while the simple interest rate on the second loan was 4.7%. At the end of the first year, Tina paid a combined interest payment of $14,761.81. What were the amounts of the two loans?

OpenStudy (anonymous):

I started with x + y= 318,338 and 0.024x + 0.047 = 14,761.81 and now I my brain stopped

OpenStudy (anonymous):

do I use elimination or substitution method? or does it matter? am I headed the right way?

OpenStudy (mertsj):

Try it this way: x dollars borrowed at 2.4% and the rest of the money (318338-x) was borrowed at 4.7% Interest on the 2.4% money: .024x Interest on the 4.7% money: .047(318338-x) Total interest for 1 year = .024x + .047(318338-x)=14761.81

OpenStudy (mertsj):

Solve that equation

OpenStudy (anonymous):

ahhhh ok I try thank you!

OpenStudy (mertsj):

yw

hero (hero):

you wrote: x + y= 318,338 and 0.024x + 0.047 = 14,761.81 but you forgot to add y to .047, so in actuality: x + y = 318, 338 .024x + .047y = 14,761.81

OpenStudy (anonymous):

ok ty Hero, I am still stuck

OpenStudy (anonymous):

I cannot figure out how to solve this

OpenStudy (anonymous):

ok thank you I keep trying:)

OpenStudy (mertsj):

\[.024x +14961.886-.047x=14761.81\] \[-.023x=-200.076\] \[.\]\[x=8698.96=\] 2.4% money

hero (hero):

I'm going to do it again on paper just to make sure my solution is correct

OpenStudy (mertsj):

It is off by 188.71

hero (hero):

yeah, I realize that now

hero (hero):

I wrote down the equation wrong again

OpenStudy (mertsj):

You made a typo when you typed the total amount of interest. It is 14761.81 NOT 13761.81

hero (hero):

I know that Mertsj. I just said that. Repeating what I already stated doesn't help.

OpenStudy (mertsj):

Sorry

hero (hero):

Just forget about my solution.